Carolyn Ellis is setting up an annuity for her retirement. She can set aside $2000 at the end of each year for the next 20 years and it will earn 6% annual interest. What lump sum will she need to set aside today at 6% annual interest to have the same retirement fund available 20 years from now? How much more will Carolyn need to invest in periodic payments than she will if she makes a lump sum payment if she intends to accumulate the same retirement balance?
Answer by rfer(10430)

FV=Pmt((1+r)^n-1/r)
FV=2000((1+.06)^20-1/.06)
FV=$73566.67
--------------------
P=FV/(1+r)n
P=73566.67/(1+.06)^20
P=$22940.00
---------------------
40000-22940=$17060.00 more with payments.